The length of the shortest side of a right triangle is 6 inches. The lengths of the other two sides are represented by consecutive odd integers. What is the lengths of the other two sides?
Answer by stanbon(75887)

The length of the shortest side of a right triangle is 6 inches. The lengths of the other two sides are represented by consecutive odd integers. What is the lengths of the other two sides?
-----------------------------------
shortest: 6 inches
hypotenuse: 2x+1
other: 2x-1
--------
Pythagoras:
6^2 + (2x-1)^2 = (2x+1)^2
-----------------------------
36 + 4x^2 -4x + 1 = 4x^2 + 4x + 1
------------
8x = 36
x = 9/2
------
hypotenuse = 2x+1 = 9+1 = 10 inches
other = 2x-1 = 8 inches
----------------------------
Cheers,
Stan H.